# leftmargin=*

### v3

@@ IV-E SEIKO Controller
We assume that actual joint positions under flexibility cannot be directly measured but can be estimated from the model. Despite model errors, our approach relies on the model’s derivatives direction to provide sufficient information about system evolution. The controller uses differentiation of the equilibrium equation with flexibility (5) to model how contact wrench distribution changes with joint command changes $\Delta\bm{\theta}^{\text{cmd}}$. This approach generalizes previously used admittance control laws such as “foot difference control” [21] which implicitly depends on flexibility without considering it.
A unique feedback law is applied from measured wrenches:
@@
$\left\lVert\mathsf{FK}_{\text{off}}(\bm{q}^{\text{flex}})\oplus\bm{J}_{\text{off}}(\bm{q}^{\text{flex}})\Delta\bm{q}^{\text{flex}}\ominus\bm{X}_{\text{off}}^{\text{d}}\right\rVert^{2}$ (11c)
@@
$\left\lVert\Delta\bm{\theta}^{\text{cmd}}\right\rVert^{2}+$ (11e)
@@
$\bm{T}_{\text{B}}\begin{bmatrix}\Delta\bm{q}^{\text{flex}}\\ \Delta\bm{\lambda}^{\text{flex}}\end{bmatrix}+\bm{t}_{\text{B}}=\bm{0}$ (11h)
$\mathsf{FK}_{\text{on}}(\bm{q}^{\text{flex}})\oplus\bm{J}_{\text{on}}(\bm{q}^{\text{flex}})\Delta\bm{q}^{\text{flex}}\ominus\bm{X}^{\text{target}}_{\text{on}}=\bm{0}$ (11i)
$\bm{\theta}^{\text{min}}\leq\bm{\theta}^{\text{cmd}}+\Delta\bm{\theta}^{\text{cmd}}\leq\bm{\theta}^{\text{max}}$ (11j)
$-\tilde{\bm{\tau}}^{\text{max}}\leq\bm{\tau}^{\text{flex}}+\Delta\bm{\tau}^{\text{flex}}\leq\tilde{\bm{\tau}}^{\text{max}}.$ (11k)
@@
Equality constraints enforce differentiated equilibrium equation with flexibility (11h) through upper floating base rows decomposition (6) and ensure no Cartesian motion for enabled contacts (11i). Inequality constraints ensure kinematic limits of joint position commands $\bm{\theta}^{\text{cmd}}$ (11j) and restrict maximum joint torques (11k).
@@ IV-F State Estimation and Effectors Admittance
This filter enhances closed-loop stability by mitigating dynamical effects affecting $\bm{\lambda}^{\text{read}}$ neglected by the quasi-static assumption. It introduces a trade-off between the reactive measurement and the term estimated through the integration of the predicted change $\Delta\bm{\lambda}^{\text{flex}}$.
@@
where filtering involves a deadband and output clamping. This control law aims to reduce interaction wrenches to zero for disabled effectors, preventing large unintended and unmodeled forces during contact establishment, facilitating foot alignment with surface orientation, and minimizing residual wrenches after contact removal. Implemented at input of the Retargeting level, this approach seamlessly integrates with operator command processing (8).
@@ V-A Implementation Details
We implemented SEIKO in C++ using RBDL [34] and Pinocchio [35] rigid body libraries. More specifically, Pinocchio efficiently computes the analytical derivatives of the terms appearing in the differentiated equation (2). We solve the QP problems using the QuadProg [36] solver.
@@
The Talos robot, manufactured by PAL Robotics, is a humanoid robot of $1.75$ m height with 32 DoFs. Externally, we measured its actual total mass to be $99.7$ kg, while the URDF model provided by PAL assumes a mass of $93.4$ kg. This discrepancy of $6$ kg can be seen by the Force-Torque sensors in the feet, which enable our controller to adapt to this model error. We changed the robot’s right hand and forearm with a 3D printed part that replaced the gripper and wrist joints beyond the elbow joint. The ball-shaped hand allows us to apply high contact forces (up to $30$ kg) on the arm during multi-contact tests. After removing the right forearm joints and excluding the head joints, our QP solver works with $n=25$ joints.
Throughout all our evaluations, we employed as flexibility model $\bm{K}$ the position-control P gains imported from PAL’s Gazebo simulation of the Talos robot. Unlike other works [10] that estimate precise flexibility model, our approach does not heavily depend on model accuracy. This is because our differentiated formulation utilizes only the approximate “gradient” direction for whole-body control.
@@ V-B Wrench Distribution Tracking
It is worth noting that we didn’t perform any identification or tuning of the robot flexibility model on the actual hardware, which may have significant errors. Estimating this flexibility [10] could enhance tracking accuracy, given that we observed near-perfect tracking performance in the Gazebo simulator which uses an ideal model.
The attached video [^2] https://hucebot.github.io/seiko_controller_website/ demonstrates additional multi-contact scenarios, such as stair climbing and stepping on a sloped surface (Fig. 1).
@@ V-D Whole Body Damping
Imperfect stiff position control and flexibilities lead to small oscillations when disturbed, particularly noticeable on Talos in the sagittal plane, causing forward-backward oscillations. In equation (10), the controller’s feedback law employs a damping term with the gain parameter $K_{d}$. We show in Fig. 5 that this unique feedback law on contact wrenches effectively attenuates these whole-body oscillations.
@@
V-E Far Reaching with Model Errors
Fig. 6 illustrates the capability of our approach to perform challenging far-reaching tasks near feasibility limits, even in the presence of large model errors. We teleoperated the right hand of the Talos robot for a forward-reaching motion as far as allowed by the controller, and added a $9$ kg load during operation on the hand to induce mass model errors. The robot remained stable thanks to the tracking of foot contact wrenches and adaptation of the whole body posture. Additionally, the Controller through equation (12) prevents excessive violation of joint torques, with a limit ratio set to $\frac{|\tau^{\text{read}}|}{\tau^{\text{max}}}<0.6$.
@@ V-F Robustness Evaluation
We performed a comprehensive analysis of our approach’s robustness using the MuJoCo simulator, as summarized in Fig. 7. The focus was on evaluating the impact of model errors and motion speed on system’s balance. We simulated the Talos robot in double support, executing $10$ motion sequences reaching a distant target with the left hand and returning to the initial posture. The number of successful trials without fall for three conditions are reported: (i) without SEIKO Controller, (ii) with SEIKO Controller but without considering joint torque limits (11k), (12), and (iii) using the full control method. Variations included hand Cartesian motion velocity (slow $2$ cm/s to fast $40$ cm/s) and additional mass on the left hand (none to $12$ kg).
@@
SEIKO Retargeting without whole-body control (left) operates in open-loop and is partially robust to motion speed but struggles with model errors. Introducing SEIKO Controller (middle) significantly improves success rates, adapting joint position commands to handle additional hand mass for balance. However, unplanned posture changes and model errors near full extension reach actuator torque limits, leading to loss of control authority. Considering actuator torque limits in the controller (right) enhances robustness by optimizing posture and avoiding infeasible hand pose commands. Challenges persist at high speeds and heavy masses, where inertial effects violate the quasi-static assumption.
@@ VI Discussion and Conclusion
Our control architecture’s robustness is showcased at moderate motion speeds (Fig. 7), but it inherently relies on the quasi-static assumption and is unsuitable for highly dynamic motions. Establishing contact with stiff position-controlled robots requires precise and slow operator commands, even if effectors admittance (14) helps mitigating this problem. Future work could explore applying the proposed approach to robots using joint impedance control. As analyzed in [10], we noted greater leg flexibility in the Talos robot than in our basic model. Although our controller enables successful contact transitions in teleoperated tasks, this significant difference hampers the quick contact switches needed for walking. Refining the flexibility model may allow walking capabilities.
@@
The robot fell when attempting to climb large $20$ cm stairs due to exceeding arm joint torque limits during the challenging contact switch. Despite being theoretically feasible according to the retargeting model, the adaptation of joint torque limits (12) is insufficient to ensure robustness if an infeasible contact transition is attempted due to model errors (e.g., underestimating the robot’s weight).
@@
Our approach overcomes the inherent lack of direct control authority over contact forces of position-controlled by explicitly considering flexibilities. While torque-controlled robots are traditionally used to perform pushing and multi-contact tasks, our SEIKO control pipeline extends these capabilities to position-controlled robots. We also demonstrate robustness to model errors, safely carrying substantial unmodelled loads at arm’s length. The unified formulation employs a single feedback law on contact forces, effectively leveraging both posture change (i.e., CoM displacement) and contact force redistribution to regulate whole-body balance. Given that the primary advantage of humanoids and other multi-limbed robots lies in their strong versatility, this research paves the way for broadening the application and deployment of real-world scenarios, utilizing more capable and adaptable multi-contact systems in uncertain contexts and environments.

### v4

@@ IV-E SEIKO Controller
We assume that actual joint positions under flexibility cannot be directly measured but can be estimated from the model. Despite model errors, our approach relies on the model’s derivatives direction to provide sufficient information about system evolution. The controller uses the derivative-based linear approximation of the equilibrium equation with flexibility (5) to model how contact wrench distribution changes with joint command changes $\Delta\bm{\theta}^{\text{cmd}}$. This approach generalizes previously used admittance control laws such as “foot difference control” [8] which implicitly depends on flexibility without considering it.
The following feedback law regulates contact wrenches. It is the only feedback effect in our unified formulation that uses measured quantities and is tuned with only two parameters:
@@
$\left\lVert\mathsf{FK}_{\text{off}}(\bm{q}^{\text{flex}})\oplus\bm{J}_{\text{off}}(\bm{q}^{\text{flex}})\Delta\bm{q}^{\text{flex}}\ominus\bm{X}_{\text{off}}^{\text{d}}\right\rVert^{2}+$ (11c)
@@
$\left\lVert\Delta\bm{\theta}^{\text{cmd}}\right\rVert^{2}$ (11e)
@@
$\bm{T}_{\text{B}}\begin{bmatrix}\Delta\bm{q}^{\text{flex}}\\ \Delta\bm{\lambda}^{\text{flex}}\end{bmatrix}+\bm{t}_{\text{B}}=\bm{0}$ (11f)
$\mathsf{FK}_{\text{on}}(\bm{q}^{\text{flex}})\oplus\bm{J}_{\text{on}}(\bm{q}^{\text{flex}})\Delta\bm{q}^{\text{flex}}\ominus\bm{X}^{\text{target}}_{\text{on}}=\bm{0}$ (11g)
$\bm{\theta}^{\text{min}}\leq\bm{\theta}^{\text{cmd}}+\Delta\bm{\theta}^{\text{cmd}}\leq\bm{\theta}^{\text{max}}$ (11h)
$-\tilde{\bm{\tau}}^{\text{max}}\leq\bm{\tau}^{\text{flex}}+\Delta\bm{\tau}^{\text{flex}}\leq\tilde{\bm{\tau}}^{\text{max}}.$ (11i)
@@
Equality constraints enforce the linearized equilibrium equation with flexibility (11f) through the first upper 6 floating base rows of decomposition (6) and ensure no Cartesian motion for enabled contacts (11g). Inequality constraints ensure kinematic limits of joint position commands $\bm{\theta}^{\text{cmd}}$ (11h) and restrict maximum joint torques (11i).
@@ IV-F State Estimation and Effectors Admittance
This filter enhances closed-loop stability by mitigating dynamical effects affecting $\bm{\lambda}^{\text{read}}$ neglected by the quasi-static assumption. It introduces a trade-off between the reactive measurement and the term estimated through the integration of the predicted change $\Delta\bm{\lambda}^{\text{flex}}$. The measured contact wrench velocity $\dot{\bm{\lambda}}^{\text{read}}$ is computed using finite differences from $\bm{\lambda}^{\text{read}}$, and then it is low-pass filtered at $10$ Hz using an exponential first-order scheme.
@@
where the filtering applies a deadband and output clamping to both linear and angular vector norms, thereby rejecting peak forces and inertial effects during motion. This effect minimizes interaction wrenches for disabled effectors, reducing collision forces during contact establishment and after contact removal, while also aiding in aligning feet with surface orientation. Implemented at input of the Retargeting level, this approach seamlessly integrates with operator command processing (8).
@@ V-A Implementation Details
We implemented SEIKO in C++ using RBDL (due to historical reasons) and Pinocchio [31] rigid body libraries. More specifically, Pinocchio efficiently computes the analytical derivatives of the terms appearing in the equation (2). We solve the QP problems using the QuadProg [32] solver.
@@
The Talos robot, manufactured by PAL Robotics, is a humanoid robot of $1.75$ m height with 32 DoFs. We measured with an independent weighing scale its actual total mass to be $99.7$ kg, while the URDF model provided by PAL assumes a mass of $93.4$ kg. This discrepancy of $6$ kg can be seen by the Force-Torque sensors in the feet, which enable our controller to adapt to this model error. We changed the robot’s right hand and forearm with a 3D printed part that replaced the gripper and wrist joints beyond the elbow joint. The ball-shaped hand (point contact) allows us to apply high contact forces (up to $30$ kg) on the arm during multi-contact tests. After removing the right forearm joints and excluding the head joints, our QP solver works with $n=25$ joints. All joints are used in position-controlled mode.
Throughout all our evaluations, we used as flexibility model $\bm{K}$ the position-control P gains imported from PAL’s Gazebo simulation of the Talos robot. Unlike other work [13] that estimate precise flexibility model, our approach does not heavily depend on model accuracy. This is because our formulation with derivatives utilizes only the approximate “gradient” direction for whole-body control.
@@ V-B Wrench Distribution Tracking
We did not perform any identification or tuning of the robot flexibility model on the actual hardware, which may have significant errors. Estimating this flexibility [13] could enhance tracking accuracy, given that we observed near-perfect tracking performance in the Gazebo simulator which uses an ideal model.
The attached video [^2] Additional videos: https://hucebot.github.io/seiko_controller_website/ demonstrates additional multi-contact scenarios, such as stair climbing and stepping on sloped surfaces (Fig. 1). The observed motions of the robot are deliberately slow due to the focus on quasi-static movements. We also conducted additional comparisons with the prior method effector admittance control [17] in Section -A in supplementary material.
@@ V-D Whole Body Damping
Imperfect stiff position control and flexibilities lead to small oscillations when disturbed, particularly noticeable on Talos in the sagittal plane, causing forward-backward oscillations. In equation (10), the controller’s feedback law employs a damping term with the gain parameter $K_{d}$. In Fig. 5, we show that this feedback law applied to contact wrenches, serving as the only feedback mechanism in our formulation using sensor measurements, effectively attenuates the whole-body oscillations.
@@
V-E Far Reaching with Model Errors
Fig. 6 illustrates the capability of our approach to perform challenging far-reaching tasks near feasibility limits, even in the presence of large model errors. We teleoperated the right hand of the Talos robot for a forward-reaching motion as far as allowed by the controller, and added a $9$ kg load during operation on the hand to induce mass model errors. The robot remained stable thanks to the tracking of foot contact wrenches and adaptation of the whole-body posture. Additionally, the Controller through equation (12) prevents excessive violation of joint torques, with a limit ratio set to $\frac{|\tau^{\text{read}}|}{\tau^{\text{max}}}<0.6$.
@@ V-F Robustness Evaluation
We performed a comprehensive analysis of our approach’s robustness using the MuJoCo simulator, as summarized in Fig. 7. The focus was on evaluating the impact of model errors and motion speed on system’s balance. We simulated the Talos robot in double support, executing $10$ motion sequences reaching a distant target with the left hand and returning to the initial posture. The number of successful trials without fall for three conditions are reported: (i) without SEIKO Controller, (ii) with SEIKO Controller but without considering joint torque limits (11i), (12), and (iii) using the full control method. Variations included hand Cartesian motion velocity (slow $2$ cm/s to fast $40$ cm/s) and additional mass on the left hand (none to $12$ kg).
@@
SEIKO Retargeting without whole-body control (left) operates in open-loop and is partially robust to motion speed but struggles with model errors. Using SEIKO Controller (middle) significantly improves success rates, adapting joint position commands to handle additional hand mass for balance. However, unplanned posture adaptations and model errors near full extension reach actuator torque limits, leading to loss of control authority. Considering actuator torque limits in the controller (right) enhances robustness by optimizing posture and avoiding infeasible hand pose commands. Challenges persist at high speeds and heavy masses, where inertial effects violate the quasi-static assumption.
@@ VI Discussion and Conclusion
Our control architecture’s robustness is showcased at moderate motion speeds (Fig. 7), but it inherently relies on the quasi-static assumption and is unsuitable for highly dynamic motions. Exploring more dynamic and agile motions is an avenue for future research. Establishing contact with stiff position-controlled robots requires precise and slow operator commands, even if effectors admittance (14) helps mitigating this problem. Future work could explore applying the proposed approach to robots using joint impedance control. As analyzed in [13], we noted greater leg flexibility in the Talos robot than in our basic model. Although our controller enables successful contact transitions in teleoperated tasks, this significant difference hampers the quick contact switches needed for walking. Refining the flexibility model may allow walking capabilities.
@@
The robot fell when attempting to climb large $20$ cm stairs for exceeding arm joint torque limits during the challenging contact switch. Despite being theoretically feasible according to the retargeting model, the adaptation of joint torque limits (12) is insufficient to ensure robustness if an infeasible contact transition is attempted due to model errors (e.g., underestimating the robot’s weight).
@@
SEIKO Controller overcomes the inherent lack of direct control authority over contact forces of position-controlled by explicitly considering flexibilities. The whole-body multi-contact formulation is grounded in model and enhances robustness to moderate motion speeds and model errors, safely carrying substantial unmodelled loads at arm’s length. The unified whole-body formulation employs a single feedback law on contact forces, effectively leveraging both postural change (i.e., CoM displacement) and contact force redistribution to regulate balance. Given that the primary advantage of humanoids and other multi-limbed robots lies in their strong versatility, this research paves the way for broadening the application and deployment of real-world scenarios, utilizing more capable and adaptable multi-contact systems in uncertain contexts and environments.
Supplementary Material:
Multi-Contact Whole-Body Force Control
for Position-Controlled Robots
Quentin Rouxel, Serena Ivaldi, and Jean-Baptiste Mouret
Figure S1: Comparison of our SEIKO Controller and effector admittance [17] for tracking inconsistent references on Talos humanoid robot simulated in double support using Gazebo simulator (left). SEIKO Retargeting is used to generate a configuration where most of the robot’s weight is positioned above the right foot. At $t=1s$, the reference sent to the controller is overridden, requesting an equal weight distribution between the two feet, inconsistent with the desired posture and CoM position. Both controllers are initiated at $t=6s$ and foot normal force tracking is displayed for each controller (right). SEIKO Controller successfully tracks the overridden reference by shifting the CoM of the robot, while the effector admittance controller results in the robot falling.
Figure S2: Comparison between our SEIKO Controller and effector admittance [17] for hand multi-contact and large model errors. Talos humanoid robot is simulated in Gazebo, with a posture featuring both feet and the right hand in contact (left). Tracking of the right hand force is compared across several initial contact forces (right). In the second row, a large external vertical force ($200N$) is applied on the robot’s torso. The effector admittance scheme fails to track the reference when faced with large external disturbances.
-A Comparison With Effector Admittance Control
As detailed in Section II, prior studies focusing on position-controlled robots [9, 10, 11] often use the method introduced in [8] for regulating contact forces. These approaches named “effector admittance”, “foot force difference control”, or “damping control for limb ends” all operate on a similar principle. They employ an admittance feedback law applied to each effector of the robot, which adjusts its Cartesian pose reference based on desired and measured contact forces before being realized via Inverse Kinematics (IK).
To compare our method against this baseline, we specifically implemented the “damping control” approach described in [17], Section IV.C. We substituted the SEIKO Controller block in Fig. 2 with an IK module. Prior to the IK calculation, we apply the following admittance law to each enabled and disabled effector:
$\bm{X}^{\text{IK}}(t)=\bm{X}^{\text{d}}(t)\oplus\Delta\bm{r}(t),$ (15)
$\Delta\bm{r}(t+\Delta t)=\Delta\bm{r}(t)\oplus\Delta t\left(K_{\text{f}}(\tilde{\bm{\lambda}}^{\text{read}}-\bm{\lambda}^{\text{d}})\ominus K_{\text{s}}\Delta\bm{r}(t)\right),$ (16)
where $\bm{X}^{\text{IK}}(t)\in SE(3)$ is the corrected effector pose input sent to the IK, $\Delta\bm{r}(t)\in SE(3)$ is the computed admittance pose offset (slightly abusing the $SE(3)$ notation), and $K_{\text{f}},K_{\text{s}}\in\mathbb{R}$ are the manually tuned wrench and spring gains, respectively. It is worth noting that we did not implement for our comparison the two additional feedback effects described in [17], as detailed in Sections IV.A and B. These two effects only use position measurements to regulate the centroidal pose. However, our Talos robot is controlled in position with stiff gains, and no significant position errors can be measured with respect to the desired posture optimized by our SEIKO Retargeting (posture flexibility is unobservable on joint sensors).
From a theoretical standpoint, contact forces in rigid multi-contact scenarios are influenced by two key factors. Firstly, the distribution of forces among different contacts can be selected within the redundant contact nullspace without inducing motion. Secondly, this nullspace is defined by the posture, specifically the position of the CoM. As its formulation is not whole-body, the effector admittance approach only utilizes the first component of force redistribution, while our SEIKO Controller offers a unified formulation capable of leveraging both force redistribution and postural adjustments.
For instance, the effector admittance approach relies on the assumption that retracting the effector away from the surface will decrease the contact force. While this assumption holds true in many scenarios, it can lead to failure when confronted with significant model errors, flexibility, or challenging postures that bring the robot’s configuration close to its feasibility limits. The following simulation experiments highlight two cases where the effector admittance method fails to accurately track the force reference due to its inability to account for whole-body postural adjustments.
-A1 Inconsistent Reference Input
Fig. S1 presents a clear example of a simulated double support scenario where retracting the foot position away from the surface fails to decrease the contact force. This situation arises when the planner module (in this case, SEIKO Retargeting) outputs an erroneous desired reference, subsequently sent to the underlying controller. Due to the inconsistency between the desired contact force and the desired posture, the controller is unable to track both simultaneously. The effector admittance control attempts to lift the right foot, which bears most of the robot’s weight, leading to the robot’s fall. In contrast, our SEIKO Controller prioritizes balance by focusing on tracking the contact forces and adjusting the posture through the whole-body formulation.
-A2 Large Model Errors
In numerous instances where the robot’s configuration remained away from the feasibility limits, we observed that the effector admittance method effectively regulated the contact forces. However, the robustness of the controller is challenged in scenarios with large model errors. Fig. S2 illustrates a failure case where a substantial external force applied to the torso necessitates the adaptation of the robot’s posture to accurately track the desired contact forces, a capability only exhibited by the SEIKO Controller.
-B Derivative of Equilibrium Equations
In the quasi-static case, the whole-body equilibrium equation in joint space is expressed as:
$\bm{g}(\bm{q})=\bm{S}\bm{\tau}+\bm{J}(\bm{q})^{\mathsf{T}}\bm{\lambda}.$ (17)
Since this equation is nonlinear with respect to $\bm{q}$, we linearize it by approximating its derivative through consideration of small variations in the configuration $\left(\bm{q}+\Delta\bm{q},\bm{\lambda}+\Delta\bm{\lambda},\bm{\tau}+\Delta\bm{\tau}\right)$:
$\bm{g}(\bm{q}+\Delta\bm{q})=\bm{S}(\bm{\tau}+\Delta\bm{\tau})+\bm{J}(\bm{q}+\Delta\bm{q})^{\mathsf{T}}(\bm{\lambda}+\Delta\bm{\lambda}),$ (18)
which is linearized and approximated as following considering partial derivatives of the gravity vector and the contact Jacobian matrix:
$\bm{g}(\bm{q})+\frac{\partial\bm{g}}{\partial\bm{q}}(\bm{q})\Delta\bm{q}=$ $\bm{S}\bm{\tau}+\bm{S}\Delta\bm{\tau}+\bm{J}(\bm{q})^{\mathsf{T}}\bm{\lambda}+\bm{J}(\bm{q})^{\mathsf{T}}\Delta\bm{\lambda}$ (19)
$+\left(\frac{\partial\bm{J}}{\partial\bm{q}}^{\mathsf{T}}(\bm{q})\bm{\lambda}\right)\Delta\bm{q}+\left(\frac{\partial\bm{J}}{\partial\bm{q}}^{\mathsf{T}}(\bm{q})\Delta\bm{\lambda}\right)\Delta\bm{q}.$
We further neglect the following term, considering it to be of second order:
$\left(\frac{\partial\bm{J}}{\partial\bm{q}}^{\mathsf{T}}\Delta\bm{\lambda}\right)\Delta\bm{q}\approx\bm{0},$ (20)
which leads to the derivative-based linear approximation of the equilibrium equation:
$\bm{g}(\bm{q})+\frac{\partial\bm{g}}{\partial\bm{q}}(\bm{q})\Delta\bm{q}=$ $\bm{S}\bm{\tau}+\bm{S}\Delta\bm{\tau}+\bm{J}(\bm{q})^{\mathsf{T}}\bm{\lambda}+\bm{J}(\bm{q})^{\mathsf{T}}\Delta\bm{\lambda}$ (21)
$+\left(\frac{\partial\bm{J}}{\partial\bm{q}}^{\mathsf{T}}(\bm{q})\bm{\lambda}\right)\Delta\bm{q}.$
Note that if we choose instead to linearize the equilibrium equation by only considering small variations in the posture $\left(\bm{q}+\Delta\bm{q},\bm{\lambda},\bm{\tau}\right)$, then the following term appears:
$\left(\frac{\partial\bm{J}}{\partial\bm{q}}^{T}(\bm{q})\bm{\lambda}\right)\Delta\bm{q},$ (22)
which is of first order but also bilinear with respect to the decision variables $\left(\Delta\bm{q},\bm{\lambda}\right)$. However, bilinear terms cannot be expressed in a QP formulation. Therefore, instead of neglecting the term (22), which is of first order, the former expression (21) provides a better approximation of the equilibrium equation by considering $\left(\Delta\bm{q},\Delta\bm{\lambda},\Delta\bm{\tau}\right)$ as the decision variables.
In the formulation of SEIKO Retargeting, the equality constraint (9g) is obtained by setting $\bm{q}=\bm{q}^{\text{d}}$ and $\bm{\lambda}=\bm{\lambda}^{\text{d}}$ in (21).
In the formulation of SEIKO Controller, the derivative of the joint flexibility model (4) is combined with (21) by setting $\bm{q}=\bm{q}^{\text{flex}},\bm{\lambda}=\bm{\lambda}^{\text{flex}},\bm{\tau}=\bm{\tau}^{\text{flex}}$. This results in the derivative-based linear approximation of the equilibrium under flexibility:
$\bm{g}(\bm{q}^{\text{flex}})+\frac{\partial\bm{g}}{\partial\bm{q}}(\bm{q}^{\text{flex}})\Delta\bm{q}^{\text{flex}}=\leavevmode\nobreak\ \bm{S}\bm{\tau}^{\text{flex}}$ (23)
$+\bm{S}\bm{K}\Delta\bm{\theta}^{\text{cmd}}-\bm{S}\bm{K}\bm{S^{\prime}}\Delta\bm{q}^{\text{flex}}$
$+\bm{J}(\bm{q}^{\text{flex}})^{\mathsf{T}}\bm{\lambda}^{\text{flex}}+\bm{J}(\bm{q}^{\text{flex}})^{\mathsf{T}}\Delta\bm{\lambda}^{\text{flex}}$
$+\left(\frac{\partial\bm{J}}{\partial\bm{q}}^{\mathsf{T}}(\bm{q}^{\text{flex}})\bm{\lambda}^{\text{flex}}\right)\Delta\bm{q}^{\text{flex}},$
yielding equation (5).
-C Whole-Body Balance Conditions
[…]
But these conditions do not consider the whole-body state and thus overlook joint kinematic and actuator torque limits. Instead, our formulation considers the complete whole-body configuration in quasi-static equilibrium. In this case, considering the CoM is not necessary, and both our SEIKO Retargeting and Controller formulation deliberately never explicitly include CoM consideration.
A sufficient condition for the robot to be balanced is that all contacts remain asymptotically stable, i.e., in contact with the environment without any relative motion between the effector in contact and the surface. For these conditions to be satisfied, both in static and dynamic cases, it is sufficient that for each contact, the contact wrench remains within a 6d or 3d convex polytope (see Section -D). These conditions are explicitly enforced in SEIKO Retargeting QP by the inequality constraints in equation (9k).
Given our quasi-static assumption and instantaneous formulation (without considering the future), enforcing conditions at the next time step is sufficient to meet the asymptotic requirement. If these conditions hold true at the next time step, they should theoretically extend into the future. Furthermore, the quasi-static equilibrium assumption is enforced by ensuring that the computed posture and contact forces satisfy the equilibrium equation (1) through its linear approximation in equations (2) and (5).
-D Contact Stability Conditions
[…]
$f_{z}\geq 0,$ (24)
$\left|\frac{f_{x}}{f_{z}}\right|\leq\mu,\leavevmode\nobreak\ \leavevmode\nobreak\ \left|\frac{f_{y}}{f_{z}}\right|\leq\mu,$ (25)
$\left|\frac{\tau_{y}}{f_{z}}\right|\leq CoP_{\text{x}}^{\text{max}},\leavevmode\nobreak\ \leavevmode\nobreak\ \left|\frac{\tau_{x}}{f_{z}}\right|\leq CoP_{\text{y}}^{\text{max}},$ (26)
$\tau_{z}^{\text{min}}\leq\tau_{z}\leq\tau_{z}^{\text{max}},$ (27)
[…]
Additionally, a bound on the maximum normal force $f^{\text{max}}\in\mathbb{R}$ can be included to prevent excessive force from being applied, particularly on a fragile effector:
$f_{z}\leq f^{\text{max}}.$ (28)
[…]
$\bm{C}_{\text{contact}}\bm{\lambda}+\bm{c}_{\text{contact}}\geq\bm{0},$ (29)
where $\bm{C}_{\text{contact}}\in\mathbb{R}^{18\times 6}$ and $\bm{c}_{\text{contact}}\in\mathbb{R}^{18}$ for a plane contact, and $\bm{C}_{\text{contact}}\in\mathbb{R}^{6\times 3}$ and $\bm{c}_{\text{contact}}\in\mathbb{R}^{6}$ for a point contact.
In (9k), all the contacts are stacked and $\bm{\lambda}$ is replaced with $\bm{\lambda}^{\text{d}}+\Delta\bm{\lambda}^{\text{d}}$.
-E Improved Contact Switching Procedure
We enhanced the contact switching procedure compared to our prior SEIKO Retargeting work [22]. To remove a contact, the desired contact force must smoothly be reduced to zero, often necessitating a smooth adjustment of the whole-body posture. To address this, SEIKO Retargeting employs a strategy of increasing the weight associated with the regularization task (9e): by heavily penalizing the contact force, the whole-body optimization naturally diminishes the force and adjusts the posture accordingly.
In [22], we empirically observed that smoothly increasing the weight of task (9e) using exponential growth yielded satisfactory switching motions. The contact is removed from the formulation when the desired force falls below a small threshold. However, the exponential choice and tuning of this procedure were arbitrary and not based on the robot’s physical limits.
In this letter, we propose a refinement of this procedure. Instead of gradually increasing the weight of task (9e), we set it instantaneously to a high value (typically $41e$ 4). The smoothness of the transition is ensured by introducing two inequality constraints on the maximum rates of change for joint positions $\dot{\bm{\theta}}^{\text{max}}$ (9l) and contact wrenches $\dot{\bm{\lambda}}^{\text{max}}$ (9m).
While the overall generated motions between the two procedures are very similar, this updated approach requires fewer arbitrary choices and tuning. Additionally, it offers the added benefit of explicit control over transition duration by setting the rate of change limits, which has clear physical semantics.
[…]